# Trapping of electromagnetic waves by two dimensional plasma-metamaterial composite structures

### v1

@@ II The model
Note EPOCH maths parser uses $lt$ instead of $<$ and $gt$ instead of $>$ signs, see EPOCH user manual section 3.19 for details. In cases 1b or 1d the number density (given by equation(5)) is similar to case 2 (figure 2(a)), but in case 1b or 1d has $n_{y}=3$, while for case 2 we set $n_{y}=128$. Number density at time zero is plotted in case 2 only (figure 2(a)). Essentially this is a slab of plasma with top hat density of $n_{0}=10^{15}$ m-3 between 512 and 1024 grid points in $x$-direction, while on domain edges number density drops to $10^{-2}n_{0}$. For the case 2 we use number density given by equation (5). For the cases 3 and 4 the number density used is given by equation (7)
@@ III The results
In is section we present the results of our numerical simulations detailed in table 1. We also provide justification for the above number density expressions used.
One of the aims of this work is to extend previous one dimensional studies of EM wave propagation on an over-dense plasma-metamaterial [1] composite into two spatial dimensions. This is achieved in 4 numerical runs for cases 1a, 1b, 1c and 1d, see table 1 for details. EPOCH 2D was modified as explained in Appendix A. The numerical grid is essentially one dimensional: $n_{x}\times n_{y}=8192\times 3$. This is because EPOCH 2D needs at least 3 grid points in $y$-direction. 8192 number of grids in $x$-direction was chosen so that left- and right-propagating EM waves, due to periodic boundary conditions, do not collide by the end simulation time of $t_{\rm end}=100/\omega_{\rm pe}$. This is clearly seen in figure 1(a), where we show time-distance plot of $E_{z}(x,y=y_{\rm max}/2,t)/E_{0}$ when electron and ion number densities are set to $10^{-2}n_{0}$, i.e. essentially plasma is so thin, it can be regarded as a vacuum. Figure 1(a) corresponds to case 1a. In figure 1(b) we plot the same electric field component but for the case 1b, i.e. when plasma is present, but metamaterial is not. Because plasma is overdense, EM wave with frequency $\omega_{0}=\omega_{\rm pe}/1.5$, cannot propagate through plasma and a standing EM wave forms. In figure 1(c) we plot the same but for the case 1c, i.e. when plasma is absent ($n_{e,i}=10^{-2}n_{0}$), but metamaterial is present. Because metamaterial frequency is $\omega_{\rm m}=1.7\omega_{0}$, the metamaterial $\mu_{m}$ is negative so again EM wave with frequency $\omega_{0}$, cannot propagate through plasma-metamaterial region and a standing EM wave forms. In figure 1(d) we plot the same but for the case 1d, i.e. when both plasma and metamaterial are present and form a composite between grid points 512 and 1024 in $x$-direction. We now see that in the plasma metamaterial composite region which lies in the interval $25.612.8<x<$ 25.6 EM wave with frequency $\omega_{0}$ propagates with negative phase speed (and positive group speed) because effective permittivity of the plasma $\varepsilon_{p}$ and the effective permeability of the metamaterial $\mu_{m}$ are both negative. We note that our figure 1 matches closely figure 2 from Ref.[1]. Since our simulation parameters for case 1 ($\omega_{\rm pe}=1.5\omega_{0}$ and $\omega_{\rm m}=1.7\omega_{0}$) are similar, to that of Ref.[1] hence the similarly of the time-distance plots is successfully achieved. In effect, figure 1 can be regarded as a successful reproduction of 1D results of Ref.[1] now with EPOCH 2D.
@@
Cases 3 and 4 aim to study trapping of EM waves by plasma-metamaterial composite barriers. Thus, in case 3 we consider a shallow slab (a barrier) where plasma number density is dropping to $0.8n_{0}$, while in case 4, number density is dropping to $0.015n_{0}$, both in the interval between 640 and 896 grid points in $x$-direction. In the normalized units this is 12.8×640/$51212.8\times 640/512<x<12.8\times 896/$ 512 i.e. $22.416<x<$ 22.4, which can be seen as light-green color strip in the middle of figure 3(a) and a white color strip in the middle of figure 4(a). In case 3 ($(1-{\rm F})\times 1.5=1.2$) plasma metamaterial composite should be transparent to EM wave with negative phase speed propagation, while in case 4 ($(1-{\rm F})\times 1.5=0.015$) plasma metamaterial composite should be opaque to the EM wave. That is exactly what we see in figures 3(c) and 4(c): in figure 3(c) the EM wave propagates freely though the entire plasma metamaterial composite ($25.612.8<x<$ 25.6), while in figure 4(c) EM wave is blocked by the deep density drop $22.416<x<$ 22.4, hence a standing EM wave forms near the $x=16$ edge. This way, we achieve trapping of EM wave by dropping density in the plasma metamaterial composite to an appropriate value, such that $(1-{\rm F})\times 1.5<1$ condition is met. In figure 3(b) we see similar behaviour of electric field as in 2(b), but figure 4(b) is significantly different, we gather that $E_{z}(x,y=y_{\rm max}/2,t=t_{\rm end})$ amplitude is significantly reduced and the frequency of the EM wave is increased. We conjecture that only non-linear harmonics can tunnel through the density drop which acts as a barrier to negative phase speed propagation seen in figures 2(b,c) figure 3(b,c). For consistency we also check spatial structure of EM wave $E_{z}(x,y,t=t_{\rm end})/E_{0}$ in figures 3(d) and 4(d). Indeed we see in figure 4(d) that EM wave that propagates in the positive $x$-direction, does not move beyond $x>16$, as it is effectively trapped.
@@
Cases 5 and 6 also aim to study trapping of EM waves by plasma-metamaterial composite, but instead of a slab we consider two-dimensional density rectangular depletions (DRD) that can be seen in figures 5(a) and 6(a), as light-green and white color boxes. In them, the number density drops to $0.8n_{0}$ in case 5 (as in case 3) and $0.015n_{0}$ in case 6 (as in case 4). Their dimensions are $x_{\rm rect}=22-16=6$ and $y_{\rm rect}=0.64\approx 1$ i.e. $y_{\rm rect}\approx c/\omega_{0}$. The gaps between DRDs are 0.32, i.e. $\approx(1/3)\times(c/\omega_{0})$. Essentially, we wanted to add some structuring in $y$-direction with structures on the scale of EM wave inertial length of $\approx c/\omega_{0}$. The results of the last two entries from table 1 are shown in figures 5 and 6. We gather from figures 5 and 6 that shallow DRDs are transparent to EM wave, while deep ones block EM wave, as expected. Next, we focus our attention on the differences from cases 3 and 4, presented earlier in figures 3 and 4. It can be seen in figure 6(b) that electric field from has complex beating pattern (multiple harmonic beating). This is in contrast to figure 4(b) where high frequency harmonics show a regular spatial pattern (no beats). Figure 6(d) does not show any structuring in $y$-direction, which seams somewhat counter intuitive. As, for example, it seems not clear at first sight, why in-between DRDs across $y$-coordinate where normalized plasma number density is unity, i.e. in the four light-brown color strips on top and bottom of white rectangles in figure 6(a), no clear EM wave propagation seen. Our explanation is that shortness of the considered gaps, $\approx(1/3)\times(c/\omega_{0})$, is not sufficient for apparent EM wave front propagation as the wave interacts with plasma on circa $>c/\omega_{0}$ scale. Nonetheless, the beats are seen in 6(b), which means there is some propagation in the gaps, hence the beat pattern. Ideally, in retrospect, we surmise that a bigger than $n_{y}=128$ could have been considered. Nonetheless, we still think the considered $y$-coordinate structuring in cases 5 and 6 provide a useful information. Future, a more elaborate study, could potentially consider different shapes of DRDs and different separation distances, which could be a subject to next work(s).
@@
IV Summary and conclusions
The aim of this work is two-fold: (i) to extend previous 1D studies [1] of electromagnetic (EM) waves propagation in an over-dense plasma-metamaterial composite into 2D spatial dimensions and (ii) to study trapping of EM waves by the composite 2D structures that act like barriers, which is analogous to EM wave trapping by preformed density cavities in near-critical density plasmas [29]. Our study comprises of several logical steps:
@@ IV Summary and conclusions
4. Next, we wanted to study EM wave trapping by a slab of density depletion (cases 3 and 4). Thus, we set plasma number density according to equation (7). We find that in case 3 ($(1-{\rm F})\times 1.5=1.2$) plasma metamaterial composite is transparent to EM wave. In this case EM wave propagates with a negative phase speed, while in case 4 ($(1-{\rm F})\times 1.5=0.015$) plasma metamaterial composite is opaque to the EM wave. Thus, forming a standing wave at the edge of an opaque region. Here the generic conclusions are as following: (i) if $(1-{\rm F})\times n_{0}$ is the minimal number density of the plasma slab and (ii) $\rm f=\omega_{\rm pe}/\omega_{0}$ is the factor by which EM wave’s angular frequency $\omega_{0}$ is less than electron plasma frequency $\omega_{\rm pe}$, then the condition for EM wave trapping by the slab is
@@
5. Further, we aimed to study EM wave trapping by two-dimensional DRDs (cases 5 and 6). Therefore plasma number density was set according to equation (8). We find that for the sizes and shapes of DRDs considered, behaviour is similar to the plasma slab density depletion. The main difference is that electric field from. It now has complex beating pattern (figure 6(b)). Such behaviour is in contrast to figure 4(b), where where high frequency harmonics show a regular spatial pattern and no beats are present. For two-dimensional DRDs the EM wave trapping condition is also given by equation (9).
@@
In summary, the barriers for EM wave propagation are created when metamaterial spatially co-exists with a plasma density depletion in a form of a slab or two-dimensional DRDs. Such approach is analogous to EM wave trapping by preformed density cavities in near-critical density plasmas investigated by Ref. [29]. We have shown that plasma metamaterial composite allows to trap EM waves by both a slab (figure 4(c)) and DRD configurations (figure 6(c)). In both cases a standing wave at the edge of an opaque region is formed. The EM wave trapping condition is established by equation (9), which is supported by our PIC numerical simulations. In our approach metamaterial acts like an agent that lets EM wave to enter over-dense plasma. The subsequent density drop makes plasma opaque to the EM wave, effectively trapping it in form of a standing wave at the edge of the barrier (density deplition slab or two dimensional DRDs).
@@
Ref. [29] established that in a situation when a near-critical density plasma contains a preformed density cavity, EM wave is trapped and becomes a standing wave. Their 2D PIC simulations have shown that within $100-200/\omega_{0}$ the standing wave damps away. We only show numerical results with end simulation time of $100/\omega_{0}$, but confirm simular result to Ref. [29] that within few hundred $1/\omega_{0}$ EM wave indeed damps. We suggest that the wave damping can offer applications such as heat deposition or substrate materials (micro)machining [32] depending on EM wave intensity. Low intesity EM waves would simply heat the plasma, potentially, high intesity EM waves can be used for substrate materials (micro)machining. From figure 4(c) (density depletion slab case) and figure 6(c) (two-dimensional DRDs case) we see that length scale of the standing wave pathern is $(1-2)c/\omega_{0}=0.25-0.5$ m. Obviously this is the value for $n_{0}=10^{15}$ particles per m-3 used, because number density sets $\omega_{\rm pe}=1.784\times 10^{9}$ Hz rad, which, in turn, sets $\omega_{0}=\omega_{\rm pe}/1.5=1.189\times 10^{9}$ Hz rad. Setting higher number density say $n_{0}=10^{19}$ m-3, can make $c/\omega_{0}$ as small as 2.5 mm. Our results potentially suggest yet unexplored applications such as: more efficient plasma vapour deposition (because the reaction rate is proportional to the product of number densities of the reactant species), controlling EM wave propagation (EM wave trapping) in invisibility cloaks and similar.
@@
Data Availability
The data and numerical codes that support the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@ III The results
Here the grid points number 512 and 1024 in x-direction correspond to $x=12.8$ and $x=25.6$ respectively. Note EPOCH maths parser uses $lt$ instead of $<$ and $gt$ instead of $>$ signs, see EPOCH user manual section 3.19 for details. In cases 1b or 1d the number density (given by equation(5)) is similar to case 2 (figure 2(a)), but in case 1b or 1d has $n_{y}=3$, while for case 2 we set $n_{y}=128$. Number density at time zero is plotted in case 2 only (figure 2(a)). Essentially this is a slab of plasma with top hat density of $n_{0}=10^{15}$ m-3 between 512 and 1024 grid points in $x$-direction, while on domain edges number density drops to $10^{-2}n_{0}$. For the case 2 we use number density given by equation (5). For the cases 3 and 4 the number density used is given by equation (7)
@@
III.1 Quasi 1D Results
One of the aims of this work is to extend previous one dimensional studies of EM wave propagation on an over-dense plasma-metamaterial [1] composite into two spatial dimensions. This is achieved in 4 numerical runs for cases 1a, 1b, 1c and 1d, see table 1 for details. The cases 1(a) to 1(d) are considered in order to validate EPOCH 2D by comparing its results with Ref. [1]. EPOCH 2D was modified as explained in Appendix A. The numerical grid is essentially one dimensional: $n_{x}\times n_{y}=8192\times 3$. This is because EPOCH 2D needs at least 3 grid points in $y$-direction. 8192 number of grids in $x$-direction was chosen so that left- and right-propagating EM waves, due to periodic boundary conditions, do not collide by the end simulation time of $t_{\rm end}=100/\omega_{\rm pe}$. This is clearly seen in figure 1(a), where we show time-distance plot of $E_{z}(x,y=y_{\rm max}/2,t)/E_{0}$ when electron and ion number densities are set to $10^{-2}n_{0}$, i.e. essentially plasma is so rarefied (well below under critical density), it can be regarded as a vacuum. Figure 1(a) corresponds to case 1a. In figure 1(b) we plot the same electric field component but for the case 1b, i.e. when plasma is present, but metamaterial is not. Because plasma is overdense, EM wave with frequency $\omega_{0}=\omega_{\rm pe}/1.5$, cannot propagate through plasma and a standing EM wave forms. In figure 1(c) we plot the same but for the case 1c, i.e. when plasma is absent ($n_{e,i}=10^{-2}n_{0}$), but metamaterial is present. Because metamaterial frequency is $\omega_{\rm m}=1.7\omega_{0}$, the metamaterial $\mu_{m}$ is negative so again EM wave with frequency $\omega_{0}$, cannot propagate through plasma-metamaterial region and a standing EM wave forms. In figure 1(d) we plot the same but for the case 1d, i.e. when both plasma and metamaterial are present and form a composite between grid points 512 and 1024 in $x$-direction. We now see that in the plasma metamaterial composite region which lies in the interval $25.612.8<x<$ 25.6 EM wave with frequency $\omega_{0}$ propagates with negative phase speed (and positive group speed) because effective permittivity of the plasma $\varepsilon_{p}$ and the effective permeability of the metamaterial $\mu_{m}$ are both negative. We note that our figure 1 matches closely figure 2 from Ref.[1]. Since our simulation parameters for case 1 ($\omega_{\rm pe}=1.5\omega_{0}$ which corresponds to $\varepsilon_{p}=-1.25$ and $\omega_{\rm m}=1.7\omega_{0}$ which corresponds to $\mu_{m}=-1.89$) are similar, to that of Ref.[1] hence the similarly of the time-distance plots is successfully achieved.
In conclusion, figure 1 is a successful reproduction of 1D results of Ref.[1], but now with EPOCH 2D.
@@ III.1 Quasi 1D Results
Cases 3 and 4 aim to study blocking of EM waves by plasma-metamaterial composite barriers. Thus, in case 3 we consider a shallow slab (a barrier) where plasma number density is dropping to $0.8n_{0}$, while in case 4, number density is dropping to $0.015n_{0}$, both in the interval between 640 and 896 grid points in $x$-direction. In the normalized units this is 12.8×640/$51212.8\times 640/512<x<12.8\times 896/$ 512 i.e. $22.416<x<$ 22.4, which can be seen as light-green color strip in the middle of figure 3(a) and a white color strip in the middle of figure 4(a). In case 3 ($(1-{\rm F})\times 1.5=1.2$) plasma metamaterial composite should be transparent to EM wave with negative phase speed propagation, while in case 4 ($(1-{\rm F})\times 1.5=0.015$) plasma metamaterial composite should be opaque to the EM wave. That is exactly what we see in figures 3(c) and 4(c): in figure 3(c) the EM wave propagates freely though the entire plasma metamaterial composite ($25.612.8<x<$ 25.6), while in figure 4(c) EM wave is blocked by the deep density drop $22.416<x<$ 22.4, hence a standing EM wave forms near the $x=16$ edge. This way, we achieve blocking of EM wave by dropping density in the plasma metamaterial composite to an appropriate value, such that $(1-{\rm F})\times 1.5<1$ condition is met. In figure 3(b) we see similar behaviour of electric field as in 2(b), but figure 4(b) is significantly different, we gather that $E_{z}(x,y=y_{\rm max}/2,t=t_{\rm end})$ amplitude is significantly reduced and the frequency of the EM wave is increased. We conjecture that only non-linear harmonics can tunnel through the density drop which acts as a barrier to negative phase speed propagation seen in figures 2(b,c) figure 3(b,c). For consistency we also check spatial structure of EM wave $E_{z}(x,y,t=t_{\rm end})/E_{0}$ in figures 3(d) and 4(d). Indeed we see in figure 4(d) that EM wave that propagates in the positive $x$-direction, does not move beyond $x>16$, as it is effectively blocked.
@@
Cases 5 and 6 also aim to study blocking of EM waves by plasma-metamaterial composite, but instead of a slab we consider two-dimensional density rectangular depletions (DRD) that can be seen in figures 5(a) and 6(a), as light-green and white color boxes. In them, the number density drops to $0.8n_{0}$ in case 5 (as in case 3) and $0.015n_{0}$ in case 6 (as in case 4). Their dimensions are $x_{\rm rect}=22-16=6$ and $y_{\rm rect}=0.64\approx 1$ i.e. $y_{\rm rect}\approx c/\omega_{0}$. The gaps between DRDs are 0.32, i.e. $\approx(1/3)\times(c/\omega_{0})$. Essentially, we wanted to add some structuring in $y$-direction with structures on the scale of EM wave inertial length of $\approx c/\omega_{0}$. The results of the last two entries from table 1 are shown in figures 5 and 6. We gather from figures 5 and 6 that shallow DRDs are transparent to EM wave, while deep ones block EM wave, as expected. Next, we focus our attention on the differences from cases 3 and 4, presented earlier in figures 3 and 4. It can be seen in figure 6(b) that electric field from has complex beating pattern (multiple harmonic beating). This is in contrast to figure 4(b) where high frequency harmonics show a regular spatial pattern (no beats). Figure 6(d) does not show any structuring in $y$-direction, which seams somewhat counter intuitive. As, for example, it seems not clear at first sight, why in-between DRDs across $y$-coordinate where normalized plasma number density is unity, i.e. in the four light-brown color strips on top and bottom of white rectangles in figure 6(a), no clear EM wave propagation seen. Our explanation is that shortness of the considered gaps, $\approx(1/3)\times(c/\omega_{0})$, is not sufficient for apparent EM wave front propagation as the wave interacts with plasma on circa $>c/\omega_{0}$ scale. Nonetheless, the beats are seen in 6(b), which means there is some propagation in the gaps, hence the beat pattern. Ideally, in retrospect, we surmise that a bigger than $n_{y}=128$ could have been considered. Nonetheless, we still think the considered $y$-coordinate structuring in cases 5 and 6 provide a useful information. Future, a more elaborate study, could potentially consider different shapes of DRDs and different separation distances, which could be a subject to next work(s).
@@
In conclusion, here we established that plasma-metamaterial composite can block EM waves by both slab and DRD configurations. This happens by forming a standing wave that subsequently damps at the edge of an opaque region.
IV Conclusions
The aim of this work is two-fold: (i) to extend previous 1D studies [1] of electromagnetic (EM) waves propagation in an over-dense plasma-metamaterial composite into 2D spatial dimensions and (ii) to study blocking of EM waves by the composite 2D structures that act like barriers, which is analogous to EM wave trapping by preformed density cavities in near-critical density plasmas [26]. Our study comprises of several logical steps:
@@ IV Conclusions
4. Next, we wanted to study EM wave blocking by a slab of density depletion (cases 3 and 4). Thus, we set plasma number density according to equation (7). We find that in case 3 ($(1-{\rm F})\times 1.5=1.2$) plasma metamaterial composite is transparent to EM wave. In this case EM wave propagates with a negative phase speed, while in case 4 ($(1-{\rm F})\times 1.5=0.015$) plasma metamaterial composite is opaque to the EM wave. Thus, forming a standing wave at the edge of an opaque region. Here the generic conclusions are as following: (i) if $(1-{\rm F})\times n_{0}$ is the minimal number density of the plasma slab and (ii) $\rm f=\omega_{\rm pe}/\omega_{0}$ is the factor by which EM wave’s angular frequency $\omega_{0}$ is less than electron plasma frequency $\omega_{\rm pe}$, then the condition for EM wave blocking by the slab is
@@
5. Further, we aimed to study EM wave blocking by two-dimensional DRDs (cases 5 and 6). Therefore plasma number density was set according to equation (8). We find that for the sizes and shapes of DRDs considered, behaviour is similar to the plasma slab density depletion. The main difference is that electric field from. It now has complex beating pattern (figure 6(b)). Such behaviour is in contrast to figure 4(b), where where high frequency harmonics show a regular spatial pattern and no beats are present. For two-dimensional DRDs the EM wave blocking condition is also given by equation (9).
@@
In summary, the barriers for EM wave propagation are created when metamaterial spatially co-exists with a plasma density depletion in a form of a slab or two-dimensional DRDs. Such approach is analogous to EM wave trapping by preformed density cavities in near-critical density plasmas investigated by Ref. [26]. We have shown that plasma metamaterial composite allows to block EM waves by both a slab (figure 4(c)) and DRD configurations (figure 6(c)). In both cases a standing wave at the edge of an opaque region is formed. The EM wave blocking condition is established by equation (9), which is supported by our PIC numerical simulations. In our approach metamaterial acts like an agent that lets EM wave to enter over-dense plasma. The subsequent density drop makes plasma opaque to the EM wave, effectively blocking it in form of a standing wave at the edge of the barrier (density depletion slab or two dimensional DRDs).
@@
Ref. [26] established that in a situation when a near-critical density plasma contains a preformed density cavity, EM wave is trapped and becomes a standing wave. Their 2D PIC simulations have shown that within $100-200/\omega_{0}$ the standing wave damps away. We only show numerical results with end simulation time of $100/\omega_{0}$, but confirm similar result to Ref. [26] that within few hundred $1/\omega_{0}$ EM wave indeed damps. We suggest that the wave damping can offer applications such as heat deposition or substrate materials (micro)machining [29] depending on EM wave intensity. Low intensity EM waves would simply heat the plasma, potentially, high intensity EM waves can be used for substrate materials (micro)machining. From figure 4(c) (density depletion slab case) and figure 6(c) (two-dimensional DRDs case) we see that length scale of the standing wave pattern is $(1-2)c/\omega_{0}=0.25-0.5$ m. Obviously this is the value for $n_{0}=10^{15}$ particles per m-3 used, because number density sets $\omega_{\rm pe}=1.784\times 10^{9}$ Hz rad, which, in turn, sets $\omega_{0}=\omega_{\rm pe}/1.5=1.189\times 10^{9}$ Hz rad. Setting higher number density say $n_{0}=10^{19}$ m-3, can make $c/\omega_{0}$ as small as 2.5 mm.
We would like to stress the difference in approach by Luan et al. [26] and this work. In particular, Luan et al. [26] consider situation when the strength of the electric field is high enough, electron density increases such that EM waves tends to be reflected by the plasma layer. Hence, in that case, the plasma is first allowing propagation of high-power EM pulse, and then blocking it as a function of time when electron density increases. In our case, the background number density does not change markedly as the time progresses. We do not keep it artificially static, it slowly evolves in time as PIC simulation progresses. Thus our a potential applications are different from that of Luan et al. [26]. Our results potentially suggest yet unexplored and/or unknown applications of controlling EM waves with plasma and metamaterial composite. These may include: a more efficient PVD (because the reaction rate is proportional to the product of number densities of the reactant species), controlling EM wave propagation (EM wave blocking) in invisibility cloaks and perhaps something yet to be discovered in the future.
We would like to remark that kinetic PIC electromagnetic simulations are more realistic than fluid FDTD simulations to capture the nonlinear effects in the plasma due to incident high-power EM energy. In this paper, we only consider low-power cases ($a=0.375$) where the physics remains linear. It should be interesting to study novel features provided by PIC simulations in the nonlinear case, for instance by showing the electron energy distribution modifications such as in Iwai et al. [1]. This will be subject of a subsequent work.
In the future, it would be also interesting to explore the novelties brought about by the 2D model, by studying, for example, the effect of the angle of incidence of the EM wave on the composite of plasma and metamaterial slab.
Yet another future avenue of research could be developing an analytical model to validate the numerical results. Indeed, air can be modeled as a propagating transmission line, while plasma as an evanescent one, and finally the combination of plasma with MNG material as a propagating transmission line. It would be feasible to analytically calculate their characteristic (i.e., intrinsic) impedance and propagation constant. The valuable end-product of such an analytical calculation could be reflection and transmission coefficients of the complex structure so that numerical results can be validated by such a model. Another possibility for an analytical model is to consider so-called Zakharov’s equations in order to describe the electronic plasma waves. In such model the slowly varying envelope of the electric field is coupled to the low-frequency variation of the density of the ions. Colin and Colin [30] used it in a situation when the laser enters the plasma, then part of it is backscattered through a Raman-type process and a Brillouin-type process. The Raman and Brillouin parts and the laser combine together to create an electron-plasma wave. These four waves interact in order to create a low-frequency variation of density of the ions which has itself an influence on the four preceding waves.
@@
An interesting observation can be also made as follows: When the de-Broglie wavelength of plasma particles becomes of the order of mean inter-particle distance, quantum effects such as the combined effect of Bohm potential and exchange correlation potential significantly modifies the dispersion properties of e.g. lower hybrid wave [37]. In the case when plasma density is high the exchange correlation potential can in principle dominate over the particle dispersion effects [38, 39, 40]. Broadly speaking this exchange correlation potential can be regarded as a form of the density fluctuation theory (DFT) [38]. The influence of exchange potential and Bohm force on the propagation characteristics of different type of waves in plasmas have been investigated before [38, 39, 40]. Since the number densities considered here, in principle, are high enough to invoke the quantum effects such as (i) the exchange correlation potential, (ii) Bohm potential, any future works on the subject should consider incorporating these for completeness.
Acknowledgements
The author would like to thank: (i) The two anonymous referees for useful comments that improved this manuscript and (ii) The Senior Editor Dr Francesco Taccogna for organizing useful and rigorous review process. Data availability statement. The data and numerical codes that support the findings of this study are available from the corresponding author upon reasonable request.